Skip to content
Solved Assignments

Practice Exam: Statistical Analysis, Financial Management, and Accounting Concepts

This practice exam covers statistical analysis and financial management concepts, beneficial for exam preparation.

Preview 4 of 12 pages

See the first 4 pages exactly as they appear in the document.

  1. Page 1
    Read page 1 as text

    Practice Exam: Statistical Analysis, Financial Management, and Accounting Concepts We know that for the normal distribution the Mean Β± 3 Οƒ (standard deviation ) intervals contains 99.7% of the total observations. So here, (0.07 - 3*0.01, 0.07+3*0.01) = 0.04 and 0.10 are the required numbers within which 99.7% of the students lies. Here X = BAC of students who drink five beer ~Normal(0.07,0.01), Z = Standard normal variate = 𝑋 βˆ’ 0 . 07 0 . 01 ~Normal(0,1) Required probability, P(X < 0.09) = P(Z < 0 . 09 βˆ’ 0 . 07 0 . 01 = 2 ) = 0.9773 (from standard normal table)= 97.73% Let that number be k so, P(X>k) = 0.15 οƒ° P(X < k) = 0.85 οƒ° P(Z < π‘˜ βˆ’ 0 . 07 0 . 01 ) =0.85 = P(Z < 1.0 4) ;From normal table Comparison gives, π‘˜ βˆ’ 0 . 07 0 . 01 = 1.04 οƒ° K = 0.07+1.04*0.01 = 0.0804 Thus the required number is 0.0804.

  2. Page 2
  3. Page 3
    Read page 3 as text

    In poker the cards are drawn and dealt randomly and after each game the cards are mixed up so that it is random. That is one game does not affect another game. So each hands are independent thus the correct option is b.

  4. Page 4
    Read page 4 as text

    Probability means the chance of the occurrence of that event, now according to the classical definition of probability is nothing but the number of favorable cases out of a large number of cases. So here the correct option is B. Assuming X = number of people in a family. We know that the total probability is 1. So, P( X = 7 or more) = 1 – (P(X=2) +P(X=3) +P(X=4) +P(X=5) +P(X=6)) = 1 - 0.98 = 0.02 P(X >2) = 1 - P(X =2) = 1 - 0.42 = 0.58 P(X = 3 or 4) = P(X = 3) + P(X =4) ;because if a family has 3 members it cant have 4 members thus. = 0.23+ 0.31 = 0.44

8 more pages in the full document

Unlock it once and it stays in your library, ready to chat with or turn into flashcards.

Unlock the full document
See all

Study this document with CramX

Once it is in your library, every tool can work from it.